media%2Fad5%2Fad5bf9f7-fbc4-4ada-a8b2-8e
Section Student ID Name: Part 3: Suppose that Canada increases its production of sweaters from 90 sweaters to 180 sweaters. What is the opportunity cost of producing an additional sweater? Part 4: Does Canada's production of sweaters experience constant or increasing opportunity costs? Part 5: Suppose that Canada increases its production of shirts from 90 shirts to 150 shirts. What is the opportunity cost of producing an additional shi rt? Part 6: Suppose that Canada increases its production of shirts from 150 shirts to 180 shirts. What is the opportunity cost of producing an additional shirt? Part 7: Does Canada's production of shirts experience constant or increasing opportunity costs?
Answer #1

1.

2.

Opportunity cost of producing 0 to 90 sweaters = 180-150 = 30 shirts

So, opportunity cost of additional sweater = 30/90 = .33 shirts

3.

Opportunity cost of producing 90 to 180 sweaters = 150 - 90 = 60 shirts

So, opportunity cost of additional sweater = 60/90 = .67 shirts

4.

It is a case of increasing opportunity cost as it can be observed from above two opportunity cost.

5.

Opportunity cost of producing 90 to 150 shirts = 180-90 = 90 sweaters

So, opportunity cost of additional shirts = 90/60 = 1.5 sweaters

6.

Opportunity cost of producing 150 to 180 shirts = 90 - 0  = 90 sweaters

So, opportunity cost of additional shirts = 90/30 = 3 sweaters

7.

It is a case of increasing opportunity cost as it can be observed from above two opportunity cost.